OpenStudy (anonymous):

Based on the graph below, what is the solution of the equation f(x) = g(x)? graph of function f of x equals negative x plus 0.25 and graph of function g of x equals x squared plus 3 multiplied by x plus 2 x = -3.5 and x = -0.5 x = -2 and x = -1 x = -3.5 and x = 0.25 x = -2 and x = 0.25

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (anonymous):

\[g(x) = x ^{2}+3 \times (x+2)\] \[f(x) = -x+0.25\]

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

g(x) doesn't make any sense

OpenStudy (anonymous):

Can you show me what g(x) looks like? I don't think that's right the way I wrote it

OpenStudy (anonymous):

Its similar to that graph.

OpenStudy (anonymous):

x^2+3x+2=y

OpenStudy (anonymous):

and -x+(0.25)=y